The equations of three quadratic functions are given below:


(i) y = 2x2 + 12x + 18


(ii) y = −3x2 − 5x + 3


(iii) y = x2 + 2x − 5


Determine the discriminant of each function and the number of x-intercepts that the parabola has. Choose


the correct sentence. The function described in


[1] (i) has no discriminant and the parabola has no x-intercepts


[2] (ii) has a negative discriminant and the parabola has two x-intercepts


[3] (i) has a zero discriminant and the parabola has one x-intercept


[4] (iii) has a positive discriminant and the parabola has one x-intercept


1
Expert's answer
2022-04-14T00:57:51-0400

The discriminant formula of a quadratic function y = ax2 + bx + c is D = b2 - 4ac.


"\text{(i) }y=2x^2+12x+18\\
D=12^2-4\cdot2\cdot18=0\\
\text{The parabola has one x-intercept.}"


"\text{(ii) }y=-3x^2-5x+3\\
D=(-5)^2-4\cdot(-3)\cdot3=61>0\\
\text{The parabola has two x-intercepts.}"


"\text{(iii) }y=x^2+2x-5\\
D=2^2-4\cdot1\cdot(-5)=24>0\\
\text{The parabola has two x-intercepts.}"


The correct sentence is [3] (i) has a zero discriminant and the parabola has one x-intercept.